OpenStudy (anonymous):

factor completely 8x^2+10x-8

OpenStudy (goformit100):

use Quadratic Theorem

jimthompson5910 (jim_thompson5910):

From 8x^2+10x-8, we see that a = 8, b = 10 and c = -8 Let's use the discriminant formula to see if we get perfect square or not D=b^2-4ac D=(10)^2-4(8)(-8) D=100-4(8)(-8) D=100-(-256) D=100+256 D=356

jimthompson5910 (jim_thompson5910):

Since we do not get a perfect square, we cannot factor this over the rationals.